# Holographic Integrated Data and Energy Transfer

### v1

@@ III-B4 Variables Initialization
$\lim_{|\mathcal{S}_{\mathrm{T}}|\rightarrow\infty}\frac{\int_{\mathcal{S}_{\mathrm{T}}}\left\|\mathbf{G}_{i}(\mathbf{s})\mathbf{G}^{H}_{j}(\mathbf{s})\right\|^{2}\mathrm{d}\mathbf{s}}{|\mathcal{S}_{\mathrm{T}}|}=0.$ (34)
where $\mathbf{G}_{i}$ and $\mathbf{G}_{j}$ are the channels spanning from the transmitter to two users, respectively.
@@
$E_{l}$ $=\lim_{|\mathcal{S}_{\mathrm{T}}|\rightarrow\infty}\frac{A_{\text{R}}\cos(\phi_{l})}{2Z}\left\|\bm{\psi_{l}}^{H}\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{l}(\mathbf{s})\left(\beta_{l}\mathbf{G}^{H}_{l}(\mathbf{s})\bm{\psi_{l}}+\sum_{j=1,j\neq l}^{L}\beta_{j}\mathbf{G}^{H}_{j}(\mathbf{s})\bm{\psi_{j}}+\sum_{i}\tau_{i}\mathbf{g}_{i}\right)\mathrm{d}\mathbf{s}\right\|^{2}$
@@
[…]
@@
Then, $\mathbf{j(s)}$ can be obtained by solving the equation, which is experssed as $\mathbf{j(s)}=\sum_{l=1}^{L}\beta_{l}\mathbf{G}^{H}_{l}(\mathbf{s})\bm{\psi_{l}}$, where $\beta_{l}=\sqrt{\dfrac{P_{t}}{\zeta_{l}^{2}\sum_{i=1}^{L}\frac{1}{\zeta_{i}}}}$ and $\zeta_{i}=\left|\int_{\mathcal{S}_{\mathrm{T}}}\bm{\psi}^{H}_{i}\mathbf{G}_{i}(\mathbf{s})\mathbf{G}^{H}_{i}\bm{\psi}_{i}(\mathbf{s})\mathrm{d}\mathbf{s}\right|^{2}$. ∎
@@ III-B5 Convergence and Complexity Analysis
We first briefly analyze the convergence of our proposed algorithm. According to the equivalence between $R_{\mathrm{eq}}$ and $R_{\mathrm{sum}}$, $R_{\mathrm{eq}}$ is bounded since the transmit power is limited. Besides, the updates of combiner $\bm{\Psi}^{(i+1)}$ in (18), (III-B2) and auxiliary variable $\bm{\rho}^{(i+1)})$ in (21) are monotonous, which is expressed as
$R_{\mathrm{eq}}(\bm{w}^{(i)},\bm{\Theta}^{(i)},\bm{\Psi}^{(i+1)},\bm{\rho}^{(i+1)})\leq R_{\mathrm{eq}}(\bm{w}^{(i)},\bm{\Theta}^{(i)},\bm{\Psi}^{(i)},\bm{\rho}^{(i+1)})\leq R_{\mathrm{eq}}(\bm{w}^{(i)},\bm{\Theta}^{(i)},\bm{\Psi}^{(i)},\bm{\rho}^{(i)}).$ (40)
@@
Combining (40), (41) and (42), $R_{\mathrm{eq}}$ is monotonically nonincreasing. Therefore, our proposed algorithm will achieve the convergence gradually.
@@
Then, we analyze the complexity of our proposed algorithm, which is mainly dependent on the updates of $\bm{w}$. The sampling number of the integral $\int_{\cal S_{\mathrm{T}}}$ is denoted by $I_{s}$. Generally, $I_{s}\gg K+L$ and $N_{F}\gg K+L$, since the H-IDET transmitter is nearly continuous. Firstly, the complexity for initializing $\bm{\bar{\psi}}$ and $\bm{\bar{\Theta}}$ is $\mathcal{O}(LI_{s})$, according to (35) and (36). Then, the complexity for updating $\bm{\rho}$ is $\mathcal{O}(K(K+L)I_{s})$, according to Eq. (III-B2) and Eq. (21). Besides, the complexity for updating $\bm{\psi}$ is $\mathcal{O}((K+L)^{2}I_{s})$ according to Eq. (18) and Eq. (III-B2). Moreover, $\bm{w}$ is initialized by solving (P7) with SCA, in which the convex approximation (P8) can be approximated as a second-order cone program (SOCP) problem. There are $6N_{F}$ real variables and $L+1$ SOC constraints in the SOCP problem. Specifically, one constraints have $26N_{F}+$ 2 real dimensions while the other $L$ constraints all have $2$ real dimensions. According to [35], the complexity of each turn to solve the SOCP is $\mathcal{O}((6N_{F})^{2}(6N_{F}+2L+2))$, while the number of required iterations is $\mathcal{O}(\sqrt{L+1})$. Similarly, when we update $\bm{w}$ by SCA, there are $6(K+L)N_{F}$ real variables and $L+2$ SOC constraints in the approximated SOCP problem. Specifically, two constraints have $26(K+L)N_{F}+$ 2 real dimensions while $L$ constraints have $2$ real dimensions. Then the complexity of each turn to solve the SOCP is $\mathcal{O}((6(K+L)N_{F})^{2}(12(K+L)N_{F}+2L+4))$ and the number of required iterations is $\mathcal{O}(\sqrt{L+2})$. Finally, the complexity of updating $\bm{\Theta}$ is $\mathcal{O}((K+L)I_{s}N_{F})$ according to (24). Let $I_{1}$ denote the number of iterations required by the convergence of sum-rate $R_{\mathrm{sum}}$. The overall complexity of Algorithm I is approximated as $\mathcal{O}((K+L)I_{1}I_{s}N_{F}+\sqrt{L+2}I_{1}(K+L)^{3}N_{F}^{3})$.
@@ IV-A Near-field focusing for single EU
Figure 2: Comparison with traditonal near-field channel
We characterize the advantages of EM channel over the traditional near-field channel in Fig. 2. Based on the Fresnel approximation [30], the traditional channel model is obtained under the first-order and second-order Taylor series expansion $\sqrt{1+x}=1+x/2-x^{2}/8+\mathcal{O}(x^{2})$, respectively. Observe from Fig. 2 that the EU can harvest more energy with the aid of EM channel and the gap narrows as the distance increases. This is because the traditional channel model is inaccurate outside the lower bound of the Fresnel region, while the EM channel model is still effective. The error arisen from traditional channel model gradually shrinks as the lower bound is approached.
We investigate the beam pattern of the plane formed by the user position and the linear antenna array in Fig. 3. To eliminate the effect of the distance, the harvested power at each location is normalized according to the path loss. Observe from Fig. 3 that although the EU is located at the same angle to the antenna array, the beam can focus the power at different distance. That is because the EM channel model without far-field approximation can reveal the most essential characteristics of the channel, by means of which we can obtain resolutions of both the angular and distance domain in the near field.
@@ IV-B Multi-user H-IDET system
  • $\bullet$
    Traditional fully digital (FD) MIMO: Inspired by the comparisons between H-MIMO and traditional MIMO in [22] and [23], we discretize the continuous surface of H-MIMO to represent a traditional discrete antenna array. Specifically, each discrete antenna is separated by half a wavelength, thus there are $M=\lceil\dfrac{2L_{x}}{\lambda}\rceil\times\lceil\dfrac{2L_{y}}{\lambda}\rceil$ antennas in total. Besides, the $m$-th discrete antenna occupies the region ${\mathcal{S}}_{m}=\left\{(x,y,z)\big{|}|x-s_{m,x}|^{2}+|y-s_{m,y}|^{2}\leq\frac{A_{m}}{\pi},z=0\right\}$, where $(s_{m,x},s_{m,y},0)$ is the center position and the effective aperture area is assumed to be $A_{m}=\lambda^{2}/(8\pi)$. Moreover, the beamforming for the $k$-th user at the $m$-th discrete tri-polarization antenna can be expressed as $\bm{\theta}_{k,m}(\mathbf{s})=\mathrm{rect}({\mathbf{s}}\in{\mathcal{S}}_{m})\mathbf{v}_{k,m}\in\mathbb{C}^{3},$ where $\mathrm{rect}({\mathbf{s}}\in{\mathcal{S}}_{m})=\left\{\begin{aligned} 1,\ &{\mathbf{s}}\in{\mathcal{S}}_{m}\\ 0,\ &{\mathbf{s}}\notin{\mathcal{S}}_{m}\end{aligned}\right.$ is the rectangular function and $\left\{\mathbf{v}_{k,m}\right\}_{m=1}^{M}$ is the digital beamformer for user $k$. By treating the EUs as DUs, a sum-rate maximization problem with transmit power constraints can be readily solved by the fractional programming [36], which is denoted by “FD”. Besides, a sum-rate maximization problem with energy harvesting constraints of EUs can be also solved by our framework in Algorithm I, which is denoted by “FD-IDET”.
  • $\bullet$
    Match-filtering (MF) H-MIMO: Xiong et. al. [37] proposed a “MF” scheme to maximize the desired signal for each user in an uplink H-MIMO system. We extend this scheme to the downlink case and treat DUs and EUs as the same. The beamforming vectors for the $k$-th user are set to $\bm{\theta}_{k}(\mathbf{s})=\mu_{k}\mathbf{G}^{H}_{k}(\mathbf{s})\bm{\psi}_{k},\mathbf{s}\in{\mathcal{S}}_{\mathrm{T}}$, wherein $\bm{\psi}_{k}=[0,1,0]^{T}$ under the assumption that the antennas at users are in the y-axis polarization and $\mu_{k}=\sqrt{\frac{P_{t}\int_{\mathcal{S}_{\mathrm{T}}}\left\|\mathbf{G}_{k}^{H}(\mathbf{s})\bm{\psi}\right\|^{2}\mathrm{d}\mathbf{s}}{\sum_{k=1}^{K+L}\int_{\mathcal{S}_{\mathrm{T}}}\left\|\mathbf{G}_{k}^{H}(\mathbf{s^{\prime}})\bm{\psi}\right\|^{2}\mathrm{d}\mathbf{s^{\prime}}}}$ is the normalizing factor to satisfy the power constraint.
  • $\bullet$
@@
We first characterise the impact of the transmit power versus sum-rate in Fig. 5. Observe from Fig. 5 that the sum-rate increases, when we increase the transmit power. Besides, our H-IDET scheme outperforms traditional ’FD’ and ’FD-IDET’ schemes. This is because H-IDET has a continuous aperture, which yields a larger spatial gains. Moreover, with the aid of H-IDET, the amplitude and the phase of the current over the entire plane are optimized, which can not be realized by traditional discrete MIMO. Therefore, the discrete beamforming remains unchanged in the region of each discrete antenna, which causes mismatches with the channel that varies with position. Furthermore, our IDET scheme outperforms the schemes of ’MF’ and ’FD’. This is because in ’MF’ and ’FD’ schemes, there is no efficient power allocation among EUs and DUs, which further degrades the IDET performance. Moreover, our scheme achieves the similar performance to the interference-free “Upper Bound”. By carefully designing the beamforming of H-IDET, our scheme is able to achieve efficient interference cancellation in the simulated scenario.
@@
We investigate the trade-off between the sum-rate of DUs and the energy harvesting power by EUs in Fig. 6. Observe from Fig. 6 that the sum-rate declines when EUs has a stringent energy harvesting requirement. That’s because more transmit power is allocated to EUs, which results in a reduction of sum-rate of DUs. Besides, a larger $N_{F}$ results in a higher IDET performance, since the Fourier base number $N_{F}$ represents the approximation accuracy in Eq. (25). When $N_{F}$ is not less than $(2\lfloor\frac{L_{x}}{\lambda}\rfloor+1)(2\lfloor\frac{L_{y}}{\lambda}\rfloor+1)(2\lfloor\frac{L_{z}}{\lambda}\rfloor+1)=441$, there are sufficient functional degrees of freedom for approximation [22]. Moreover, “RI-SCA” and “RI-BCD” denote the random initialization instead of our proposed scheme in the step (1) and step (5) of Algorithm I, respectively. $N_{F}$ is set to 144 in “RI-SCA” and “RI-BCD”, respectively. It is observed that the sum-rate of random initialization suddenly drops to 0 as the energy harvesting requirement of EUs increases. This is because the random initialization results in no feasible solution to the approximated convex problem. Since the initialization of $\bm{\bar{w}}$ has the most influence to iterations, “RI-SCA” drops much early than “RI-BCD”, although “RI-BCD” has a lower approximation accuracy. The robustness of our algorithm is significantly improved by our proposed initialization scheme.
@@
We characterise the impact of transmitter’s surface area $\mathrm{A_{T}}$ versus the sum-rate in Fig. 7. Observe from Fig. 7 that the sum-rate increases with a larger antenna aperture, since more spatial gains can be achieved. Besides, more DUs or less EUs both result in a higher sum-rate, sicne more EUs will compete for more transmit power while more DUs will provide a higher diversity gain for the system. Moreover, the x-intercept of the curve increases with a lager number of EUs, Since a lager antenna aperture is required to satisfy higer energy harvesting constraints.
Refer to captionRefer to captionRefer to captionRefer to caption
(a) Amplitude of H-IDET for DUs
Refer to captionRefer to captionRefer to captionRefer to caption
(b) Phase of H-IDET for DUs
Refer to captionRefer to caption
(c) Amplitude of H-IDET for EUs
Refer to captionRefer to caption
(d) Phase of H-IDET for EUs
Figure 8: The phase and amplitude of H-IDET for DUs and EUs

### v2

@@ III-B4 Variables Initialization
$\lim_{|\mathcal{S}_{\mathrm{T}}|\rightarrow\infty}\frac{\left\|\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{i}(\mathbf{s})\mathbf{G}^{H}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\right\|^{2}}{|\mathcal{S}_{\mathrm{T}}|}=0,\ \mathrm{for}\ i\neq j.$ (36)
where $\mathbf{G}_{i}$ and $\mathbf{G}_{j}$ are the channels spanning from the transmitter to two different users, respectively.
@@
$E_{l}$ $=\lim_{|\mathcal{S}_{\mathrm{T}}|\rightarrow\infty}\frac{A_{\text{R}}\cos(\phi_{l})}{2Z}\Bigg{\|}\bm{\psi_{l}}^{H}\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{l}(\mathbf{s})\Bigg{(}\beta_{l}\mathbf{G}^{H}_{l}(\mathbf{s})\bm{\psi_{l}}$
$+\sum_{j=1,j\neq l}^{L}\beta_{j}\mathbf{G}^{H}_{j}(\mathbf{s})\bm{\psi_{j}}+\sum_{i}\tau_{i}\mathbf{g}_{i}\Bigg{)}\mathrm{d}\mathbf{s}\Bigg{\|}^{2}$
@@
Therefore, the asymptotically optimal $\bm{\psi_{l}}$ for the EUl can be obtained by $\bm{\psi_{l}}=\arg\max E_{l}=\frac{\xi_{max}\left\{\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{l}(\mathbf{s})\mathbf{G}_{l}^{H}(\mathbf{s})\mathrm{d}\mathbf{s}\right\}}{\left\|\xi_{max}\left\{\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{l}(\mathbf{s})\mathbf{G}_{l}^{H}(\mathbf{s})\mathrm{d}\mathbf{s}\right\}\right\|}$. Since the second term of $\mathbf{j(s)}$ have no contribution to any user, we set $\tau_{i}=0$. Moreover, the asymptotically optimal $\mathbf{j(s)}$ is obtained when all EUs harvest the same energy. By considering the constraint of (35a), we have the following equation
@@
Then, $\mathbf{j(s)}$ can be obtained by solving the equation, which is experssed as $\mathbf{j(s)}=\sum_{l=1}^{L}\beta_{l}\mathbf{G}^{H}_{l}(\mathbf{s})\bm{\psi_{l}}$, where $\beta_{l}=\sqrt{\dfrac{P_{t}}{\zeta_{l}^{2}\sum_{i=1}^{L}\frac{1}{\zeta_{i}}}}$ and $\zeta_{i}=\left|\int_{\mathcal{S}_{\mathrm{T}}}\bm{\psi}^{H}_{i}\mathbf{G}_{i}(\mathbf{s})\mathbf{G}^{H}_{i}\bm{\psi}_{i}(\mathbf{s})\mathrm{d}\mathbf{s}\right|$. ∎
@@ III-B5 Convergence and Complexity Analysis
We first briefly analyze the convergence of our proposed algorithm. According to the equivalence between $R_{\mathrm{eq}}$ and $R_{\mathrm{sum}}$, $R_{\mathrm{eq}}$ is bounded since the transmit power is limited. Besides, the updates of combiner $\bm{\Psi}^{(i+1)}$ in (III-B2), (III-B2) and auxiliary variable $\bm{\rho}^{(i+1)})$ in (22) are monotonous, which is expressed as
$R_{\mathrm{eq}}(\bm{w}^{(i)},\bm{\Theta}^{(i)},\bm{\Psi}^{(i+1)},\bm{\rho}^{(i+1)})$ $\leq R_{\mathrm{eq}}(\bm{w}^{(i)},\bm{\Theta}^{(i)},\bm{\Psi}^{(i)},\bm{\rho}^{(i+1)})$
$\leq R_{\mathrm{eq}}(\bm{w}^{(i)},\bm{\Theta}^{(i)},\bm{\Psi}^{(i)},\bm{\rho}^{(i)}).$ (42)
@@
Combining (III-B5), (43) and (44), $R_{\mathrm{eq}}$ is monotonically nonincreasing. Therefore, our proposed algorithm will achieve the convergence gradually.
@@
Then, we analyze the complexity of our proposed algorithm, which is mainly dependent on the updates of $\bm{w}$. The sampling number of the integral $\int_{\cal S_{\mathrm{T}}}$ is denoted by $I_{s}$. Generally, $I_{s}\gg K+L$ and $N_{F}\gg K+L$, since the H-IDET transmitter is nearly continuous. Firstly, the complexity for initializing $\bm{\bar{\psi}}$ and $\bm{\bar{\Theta}}$ is $\mathcal{O}(LI_{s})$, according to (37) and (38). Then, the complexity for updating $\bm{\rho}$ is $\mathcal{O}(K(K+L)I_{s})$, according to Eq. (III-B2) and Eq. (22). Besides, the complexity for updating $\bm{\psi}$ is $\mathcal{O}((K+L)^{2}I_{s})$ according to Eq. (III-B2) and Eq. (III-B2). Moreover, $\bm{w}$ is initialized by solving (P7) with SCA, in which the convex approximation (P8) can be approximated as a second-order cone program (SOCP) problem. There are $6N_{F}$ real variables and $L+1$ SOC constraints in the SOCP problem. Specifically, one constraints have $26N_{F}+$ 2 real dimensions while the other $L$ constraints all have $2$ real dimensions. According to [44], the complexity of each turn to solve the SOCP is $\mathcal{O}((6N_{F})^{2}(6N_{F}+2L+2))$, while the number of required iterations is $\mathcal{O}(\sqrt{L+1})$. Similarly, when we update $\bm{w}$ by SCA, there are $6(K+L)N_{F}$ real variables and $L+2$ SOC constraints in the approximated SOCP problem. Specifically, two constraints have $26(K+L)N_{F}+$ 2 real dimensions while $L$ constraints have $2$ real dimensions. Then the complexity of each turn to solve the SOCP is $\mathcal{O}((6(K+L)N_{F})^{2}(12(K+L)N_{F}+2L+4))$ and the number of required iterations is $\mathcal{O}(\sqrt{L+2})$. Finally, the complexity of updating $\bm{\Theta}$ is $\mathcal{O}((K+L)I_{s}N_{F})$ according to (26). Let $I_{1}$ denote the number of iterations required by the convergence of sum-rate $R_{\mathrm{sum}}$. The overall complexity of Algorithm I is approximated as $\mathcal{O}((K+L)I_{1}I_{s}N_{F}+\sqrt{L+2}I_{1}(K+L)^{3}N_{F}^{3})$.
III-B6 Spatial orthogonality analysis
In this part, we will further discuss the spatial orthogonality in Lemma 2. For simplicity, we assume that the H-MIMO transmitter is linear along the $z$-direction, satisfying $\left\{\mathbf{s}=(0,0,s)\in\mathcal{C}|-c/2<s<c/2\right\}$. The user is located at $\mathbf{s}=(d,0,r)$. This model corresponds to single-polarized linear antennas, while the Green function of scalar radiation near-field can be derived as
$G(\mathbf{r},\mathbf{s})={\mathbf{G}}(\mathbf{r},\mathbf{s})(3,3)=\frac{\mathsf{i}\kappa Z_{0}d^{2}e^{\mathsf{i}\kappa\sqrt{(r-s)^{2}+d^{2}}}}{4\pi((r-s)^{2}+d^{2})^{\frac{3}{2}}}.$ (45)
Let two user are located at $(d_{1},0,r_{1})$ and $(d_{2},0,r_{2})$, respectively, where $d_{1}\neq d_{2}$, $r_{1}\neq r_{2}$. Then, the correlation function can be calculated as
$\left|\int_{\mathcal{C}}G^{H}(\mathbf{r_{1}},\mathbf{s})G(\mathbf{r_{2}},\mathbf{s})\mathrm{d}s\right|$
$=$ $\frac{\kappa^{2}Z_{0}^{2}d^{4}}{16\pi^{2}}\left|\int_{\mathcal{C}}\frac{e^{\mathsf{i}\kappa\left(\sqrt{(r_{1}-s)^{2}+d_{1}^{2}}-\sqrt{(r_{2}-s)^{2}+d_{2}^{2}}\right)}}{((r_{1}-s)^{2}+d_{1}^{2})^{\frac{3}{2}}((r_{2}-s)^{2}+d_{2}^{2})^{\frac{3}{2}}}\mathrm{d}s\right|.$ (46)
Figure 3: Comparison with traditonal near-field channel
We assume the users are located at the Fresnel region. The following two approximations can be adopted to phase and amplitude, respectively [38].
$\sqrt{(r-s)^{2}+d^{2}}\approx d+\frac{(r-s)^{2}}{2d},$ (47)
$(r-s)^{2}+d^{2}\approx r^{2}+d^{2}.$ (48)
Then, we have
$\left|\int_{\mathcal{C}}G^{H}(\mathbf{r_{1}},\mathbf{s})G(\mathbf{r_{2}},\mathbf{s})\mathrm{d}s\right|$
$=$ $\frac{\kappa^{2}Z_{0}^{2}d^{4}e^{\mathsf{i}\kappa\left(d_{1}-d_{2}+\frac{r_{1}^{2}}{2d_{1}}-\frac{r_{2}^{2}}{2d_{2}}\right)}}{16\pi^{2}(r_{1}^{2}+d_{1}^{2})^{\frac{3}{2}}(r_{2}^{2}+d_{2}^{2})^{\frac{3}{2}}}\left|\int_{\mathcal{C}}e^{\mathsf{i}\kappa\left(\frac{s^{2}-2r_{1}s}{2d_{1}}-\frac{s^{2}-2r_{2}s}{2d_{2}}\right)}\mathrm{d}s\right|$
$=$ $\epsilon_{0}\left|\int_{\mathcal{C}}e^{\mathsf{i}\left(\epsilon_{1}s+\epsilon_{2}s^{2}\right)}\mathrm{d}s\right|$ (49)
where $\epsilon_{0}=\frac{\kappa^{2}Z_{0}^{2}d^{4}e^{\mathsf{i}\kappa\left(d_{1}-d_{2}+\frac{r_{1}^{2}}{2d_{1}}-\frac{r_{2}^{2}}{2d_{2}}\right)}}{16\pi^{2}(r_{1}^{2}+d_{1}^{2})^{\frac{3}{2}}(r_{2}^{2}+d_{2}^{2})^{\frac{3}{2}}}$, $\epsilon_{1}=\frac{r_{2}}{d_{2}}-\frac{r_{1}}{d_{1}}$ and $\epsilon_{2}=\frac{1}{2d_{1}}-\frac{1}{2d_{2}}\neq 0$. By substituting $s+\frac{\epsilon_{1}}{2\epsilon_{2}}$ with $t$, we can obtain
$\lim_{c\rightarrow\infty}\frac{1}{c}\left|\int_{\mathcal{C}}G^{H}(\mathbf{r_{1}},\mathbf{s})G(\mathbf{r_{2}},\mathbf{s})\mathrm{d}s\right|=\lim_{c\rightarrow\infty}\frac{\epsilon_{0}}{c}\left|\int_{-\infty}^{\infty}e^{\mathsf{i}\epsilon_{2}t^{2}}\mathrm{d}t\right|.$ (50)
According to the Fresnel function [38], the limitation of the integral can be expressed as
$\int_{-\infty}^{\infty}e^{\mathsf{i}\epsilon_{2}t^{2}}\mathrm{d}t=2\int_{0}^{\infty}\left(\cos(\epsilon_{2}t^{2})+\mathsf{i}\sin(\epsilon_{2}t^{2})\right)\mathrm{d}t=\sqrt{\frac{\pi}{2\epsilon_{2}}}(1+j).$ (51)
Therefore, we can obtain
$\lim_{c\rightarrow\infty}\frac{1}{c}\left|\int_{\mathcal{C}}G^{H}(\mathbf{r_{1}},\mathbf{s})G(\mathbf{r_{2}},\mathbf{s})\mathrm{d}s\right|=\lim_{c\rightarrow\infty}\frac{1}{c}\sqrt{\frac{\pi}{\epsilon_{2}}}=0,$
$\mathrm{for}\ d_{1}\neq d_{2},r_{1}\neq r_{2},$ (52)
which completes the proof of the spatial orthogonality for single-polarized linear antennas. Similar discussions can be made for other cases such as multi-polarization and planar arrays. Besides, the speed of the convergence is affected by the locations of users. These discussions are beyond the scope of this paper and reserved for our future research.
The asymptotically EM wave received by the $k$-th DU can be derived as
$\lim_{|\mathcal{S}_{\mathrm{T}}|\rightarrow\infty}\mathbf{y}_{k}=\underbrace{x_{k}\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{k}(\mathbf{s})\mathrm{d}\mathbf{s}}_{\varrho}+o(\varrho)+\mathbf{n}_{k},$ (53)
where $f(y)\sim o(y)$ indicates that $\lim_{y\rightarrow\infty}(f(y)/y)=0$. Therefore, the inter-user interference asymptotically approach to zero and the sum-rate saturates to the interference-free sum-rate.
@@ IV-A Near-field focusing for single EU
We investigate the effect of user location on the planar beampattern in Fig. 2. The variables of each subfigure are different user locations, which are (0,0,4), (0,0,8) and (0,0,25), respectively. Therefore, the beam is also focused at different locations, respectively. To eliminate the effect of the distance, the harvested power at each location is normalized according to the path loss. In conventional far-field communication systems, the beam can only concentrate its energy at a certain angle, but not at different distances. Different from far-field beamsteering, the proposed scheme can achieve near-field focusing. Observe from Fig. 2 that although the EU is located at the same angle to the antenna array, the beam can focus the power at different distance. That is because the EM channel model without far-field approximation can reveal the most essential characteristics of the channel, by means of which we can obtain resolutions of both the angular and the distance domain in the near field.
We characterize the advantages of EM channel over the traditional near-field channel in Fig. 3. Based on the Fresnel approximation [38], the traditional channel model is obtained under the first-order and second-order Taylor series expansion $\sqrt{1+x}=1+x/2-x^{2}/8+\mathcal{O}(x^{3})$, respectively. Observe from Fig. 3 that the EU can harvest more energy with the aid of EM channel and the gap narrows as the distance increases. This is because the traditional channel model is inaccurate outside the lower bound of the Fresnel region, while the EM channel model is still effective. The error arisen from traditional channel model gradually shrinks as the lower bound is approached.
@@ IV-B Multi-user H-IDET system
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 8: Amplitude of H-IDET for DUs (a) and EUs (c). Phase of H-IDET for DUs (b) and EUs (d).
Figure 9: Near-filed focusing for the IDET system: Near-filed focusing for the IDET system: the normalized beam pattern of the whole H-MIMO (a), and the normalized beam pattern of the beamforming for EUs (b), DU 1 (c) and DU 2 (d).
  • Traditional fully digital (FD) MIMO: Inspired by the comparisons between H-MIMO and traditional MIMO in [26] and [27], we discretize the continuous surface of H-MIMO to represent a traditional discrete antenna array. Specifically, each discrete antenna is separated by half a wavelength, thus there are $M=\lceil\dfrac{2L_{x}}{\lambda}\rceil\times\lceil\dfrac{2L_{y}}{\lambda}\rceil$ antennas in total. Besides, the $m$-th discrete antenna occupies the region ${\mathcal{S}}_{m}=\left\{(x,y,z)\big{|}|x-s_{m,x}|^{2}+|y-s_{m,y}|^{2}\leq\frac{A_{m}}{\pi},z=0\right\}$, where $(s_{m,x},s_{m,y},0)$ is the center position and the effective aperture area is assumed to be $A_{m}=\lambda^{2}/(8\pi)$. Moreover, the beamforming for the $k$-th user at the $m$-th discrete tri-polarization antenna can be expressed as $\bm{\theta}_{k,m}(\mathbf{s})=\mathrm{rect}({\mathbf{s}}\in{\mathcal{S}}_{m})\mathbf{v}_{k,m}\in\mathbb{C}^{3},$ where $\mathrm{rect}({\mathbf{s}}\in{\mathcal{S}}_{m})=\left\{\begin{aligned} 1,\ &{\mathbf{s}}\in{\mathcal{S}}_{m}\\ 0,\ &{\mathbf{s}}\notin{\mathcal{S}}_{m}\end{aligned}\right.$ is the rectangular function and $\left\{\mathbf{v}_{k,m}\right\}_{m=1}^{M}$ is the digital beamformer for user $k$. By treating the EUs as DUs, a sum-rate maximization problem with transmit power constraints can be readily solved by the fractional programming [45], which is denoted by “FD”. Besides, a sum-rate maximization problem with energy harvesting constraints of EUs can be also solved by our framework in Algorithm I, which is denoted by “FD-IDET”. Both “FD-IDET” and “FD” don’t use the Fourier transform in the proposed scheme.
  • Match-filtering (MF) H-MIMO: Xiong et al. [46] proposed a “MF” scheme to maximize the desired signal for each user in an uplink H-MIMO system. We extend this scheme to the downlink case and treat DUs and EUs as the same. The beamforming vectors for the $k$-th user are set to $\bm{\theta}_{k}(\mathbf{s})=\mu_{k}\mathbf{G}^{H}_{k}(\mathbf{s})\bm{\psi}_{k},\mathbf{s}\in{\mathcal{S}}_{\mathrm{T}}$, wherein $\bm{\psi}_{k}=[0,1,0]^{T}$ under the assumption that the antennas at users are in the y-axis polarization and $\mu_{k}=\sqrt{\frac{P_{t}\int_{\mathcal{S}_{\mathrm{T}}}\left\|\mathbf{G}_{k}^{H}(\mathbf{s})\bm{\psi}\right\|^{2}\mathrm{d}\mathbf{s}}{\sum_{k=1}^{K+L}\int_{\mathcal{S}_{\mathrm{T}}}\left\|\mathbf{G}_{k}^{H}(\mathbf{s^{\prime}})\bm{\psi}\right\|^{2}\mathrm{d}\mathbf{s^{\prime}}}}$ is the normalizing factor to satisfy the power constraint.
@@
We first characterise the impact of the transmit power versus sum-rate in Fig. 5. Observe from Fig. 5 that the sum-rate increases, when we increase the transmit power. Besides, our H-IDET scheme outperforms traditional ’FD’ and ’FD-IDET’ schemes. Although a continuous aperture surface has the same DoF with discrete antenna array for the same aperture size, the performance upper bound of the system is difficult to achieved in complex scenarios with multi-user IDET and polarized electromagnetic waves. Continuous aperture surface can manipulate electromagnetic waves more flexibly over the entire plane and approximate the performance upper bound better than discrete arrays. Moreover, with the aid of H-IDET, the amplitude and the phase of the current over the entire plane are optimized, while traditional discrete MIMO can only manipulate discrete areas. Therefore, the discrete beamforming remains unchanged in the region of each discrete antenna, which causes mismatches with the channel that varies with position. Furthermore, our IDET scheme outperforms the schemes of ’MF’ and ’FD’. This is because in ’MF’ and ’FD’ schemes, there is no efficient power allocation among EUs and DUs, which further degrades the IDET performance. Moreover, our scheme achieves the similar performance to the interference-free “Upper Bound”. By carefully designing the beamforming of H-IDET, our scheme is able to achieve efficient interference cancellation in the simulated scenario.
@@
We investigate the trade-off between the sum-rate of DUs and the energy harvesting power by EUs in Fig. 6. Observe from Fig. 6 that the sum-rate declines when EUs has a stringent energy harvesting requirement. That’s because more transmit power is allocated to EUs, which results in a reduction of sum-rate of DUs. Besides, a larger $N_{F}$ results in a higher IDET performance, since the Fourier base number $N_{F}$ represents the approximation accuracy in Eq. (27). When $N_{F}$ is not less than $(2\lceil\frac{L_{x}}{\lambda}\rceil+1)(2\lceil\frac{L_{y}}{\lambda}\rceil+1)(2\lceil\frac{L_{z}}{\lambda}\rceil+1)=441$, there are sufficient functional degrees of freedom for approximation [26]. Moreover, “RI-SCA” and “RI-BCD” denote the random initialization instead of our proposed scheme in the step (1) and step (5) of Algorithm I, respectively. $N_{F}$ is set to 144 in “RI-SCA” and “RI-BCD”, respectively. It is observed that the sum-rate of random initialization suddenly drops to 0 as the energy harvesting requirement of EUs increases. This is because the random initialization results in no feasible solution to the approximated convex problem. Since the initialization of $\bm{\bar{w}}$ has the most influence to iterations, “RI-SCA” drops much early than “RI-BCD”, although “RI-BCD” has a lower approximation accuracy. The robustness of our algorithm is significantly improved by our proposed initialization scheme.
@@
We characterise the impact of transmitter’s surface area $\mathrm{A_{T}}$ versus the sum-rate in Fig. 7. Observe from Fig. 7 that the sum-rate increases with a larger antenna aperture, since more spatial gains can be achieved. Besides, more DUs or less EUs both result in a higher sum-rate, since more EUs will compete for more transmit power while more DUs will provide a higher diversity gain for the system. Moreover, the x-intercept of the curve increases with a lager number of EUs, Since a larger antenna aperture is required to satisfy higher energy harvesting constraints.